OpenStudy (karatechopper):

Find the limit. (Special Trig functions) As the lim (sinx/2x^2-x) approaches 0.

OpenStudy (karatechopper):

I broke the limit down and did this: sinx sinx ---- - ----- x 2x-1 I used the special trig function to find out that sinx/x=1. So now I am left with 1 - sinx/2x-1 And I'm stuck!

OpenStudy (karatechopper):

to make more sense- I am stuck on 1- (sinx/2x-1)

OpenStudy (irishboy123):

\[lim_{x\rightarrow 0} \frac{sin \ x}{2x^2 - x} = lim_{x\rightarrow 0} \frac{\frac{sin \ x}{x}}{2x - 1} \] plus \[lim_{x\rightarrow 0} \frac{sin}{ x} = 1\]

OpenStudy (chillout):

Still having trouble?

OpenStudy (karatechopper):

yeah I still don't understand..

OpenStudy (irishboy123):

\(\large \frac{1}{2(0) - 1} = ???\)

OpenStudy (chillout):

You need to notice that 2x²-x = x(2x-1). That's what @IrishBoy123 did.

OpenStudy (chillout):

And that's the same as \(\frac{ \sin x }{ x }*\frac{ 1 }{ 2x-1 }\)

OpenStudy (irishboy123):

the crucial thing here is that \[lim_{x\rightarrow 0} \frac{sin \ x}{x} = 1\] so you set it up to have a \( \frac{sin \ x}{x} \) term but that \( \frac{sin \ x}{x} \) term is seminal

OpenStudy (irishboy123):

this the important bit @karatechopper ! https://gyazo.com/ad5c26d3cf8e2b41c20a8c7935b34a83 splitting it out so you have a \(\frac{sinx}{x}\) term

OpenStudy (anonymous):

@ChillOut hi

OpenStudy (karatechopper):

understood! thanks :)

OpenStudy (karatechopper):

BUT. How can you just take that x to the numerator when its being multiplied in the denominator?

OpenStudy (chillout):

When you have something like \(\LARGE\frac{\frac{a}{b}}{c}\) you can rewrite that as \(\LARGE\frac{a}{b*c}\).

OpenStudy (chillout):

@karatechopper